# Towards Responsible Evaluation for Text-to-Speech

### v1

@@ 3.1 Challenges with Objective Metrics
Objective metrics are valued for scalability and reproducibility, but they face two fundamental limitations. First, the relationship between metric scores and human perception is nonlinear. Once performance exceeds a certain threshold, further improvements often bring diminishing perceptual benefits. Second, metrics derived from models embody their internal biases and uncertainty, rendering evaluation outcomes dependent not only on the input but also on the model itself.
WER and CER
To evaluate intelligibility and robustness, WER is computed by comparing transcriptions of synthetic speech from ASR systems with reference texts. While they can detect clear intelligibility failures, their reliability is limited in two aspects. First, inherent errors in ASR systems can lead to a mismatch between metric scores and actual perceptual quality, even when the synthesized speech is perceptually adequate to humans. Second, they are not linearly correlated with perceived intelligibility. At already low error rates, further score reductions have a negligible perceptual impact. As demonstrated by the experiments in Appendix C, a reduction from 1.61 to 1.47 has minimal impact on user perception.
@@
To assess speaker similarity, the SIM score is computed by the cosine similarity between speaker embeddings extracted from reference and synthesized speech. These embeddings, derived from speaker verification models like x-vectors Snyder et al. (2018) and ECAPA-TDNN Desplanques et al. (2020), can be sensitive to channel variations, background noise, and even phonetic content, leading to unstable scores. In practice, once the SIM score exceeds a certain threshold, further improvements offer limited perceptual gains. In more realistic scenarios like podcasts or audiobooks, existing metrics rarely account for speaker consistency over extended durations.
@@
Predicted MOS scores are generated by models trained on human ratings Cooper and Yamagishi (2021); Liu et al. (2025) collected following ITU-T P.808 Naderi and Cutler (2020). While these models offer a scalable alternative to human evaluation, they struggle with generalization and uncertainty estimation, primarily due to limitations in the diversity of training data and model representational power. Prior works Wang et al. (2025b); Cooper et al. (2022) have shown that existing MOS prediction models often produce inconsistent results even on in-domain data, and their performance degrades significantly when applied to out-of-domain data. A typical example of domain mismatch is the widespread use of DNSMOS Reddy et al. (2021); Cumlin et al. (2024); Reddy et al. (2022), which is trained on speech enhancement data yet commonly employed to evaluate synthesized speech. Moreover, MOS prediction models generally lack uncertainty estimation Wang et al. (2024a), as they typically provide only point estimates without associated confidence intervals, making it difficult to assess the reliability of the predicted quality scores. This remains rarely examined in current research.
@@
To assess speech prosody, most evaluation practices Galdino et al. (2025) employ log $F_{0}$ RMSE. However, this metric correlates weakly with human perceptual judgments Yang et al. (2025b), and it only captures pitch but overlooks other prosodic aspects, including rhythm, stress, and intensity Arvaniti (2020).
@@ 3.2 Challenges with Subjective Metrics
Subjective evaluation remains the primary choice for assessing perceptual quality in TTS, with MOS serving as the dominant protocol. MOS employs a five-point absolute category rating scale to rate individual utterances. Variants such as CMOS and MUSHRA are used for pairwise or comparative assessments. Although broadly regarded as the gold standard, these methods fall short in terms of sensitivity, consistency, and practical feasibility. One major drawback of MOS stems from its limited resolution. As the quality of synthetic speech continues to improve, MOS scores tend to saturate Wang et al. (2025c). This ceiling effect obscures perceptual differences between high-performing systems, making it increasingly difficult to distinguish among them. Another issue arises from the inherent variability in subjective ratings. Factors such as listener bias, contextual framing, playback conditions, and even day-to-day mood can introduce substantial noise. Without rigorous rater calibration and experimental controls, evaluations become unreliable. Moreover, the high cost associated with subjective evaluations presents a practical barrier. The process of recruiting a large and diverse pool of listeners, along with the need to ensure controlled testing conditions, demands considerable time and resources. These requirements often limit the feasibility and scale of such evaluations.
@@ 3.3 Underexplored Dimensions in TTS Evaluation
Existing evaluation dimensions in TTS fail to keep pace with the growing complexity of real-world applications. Widely used metrics capture only a narrow portion of what matters in practical synthesis scenarios. We identify three critical yet underexplored dimensions that are essential for responsible and forward-looking TTS evaluation.
@@
In real-world applications such as audiobooks and podcasts, coherence across sentences and stability in prosody and speaker identity are essential. However, most existing evaluations remain focused on short utterances. There is a lack of representative test sets and metrics specifically designed to assess long-form fluency, prosodic consistency, and discourse-level control.
@@
Recent TTS models demonstrate growing capability in synthesizing expressive speech, yet evaluation methods lag. There is no consensus on emotion taxonomies or scales for emotion intensity, and subjective metrics like emotion MOS often lack sensitivity to subtle distinctions Yang et al. (2025a). Available datasets typically contain discrete labels and limited emotional diversity.
@@
Polyphonic Word Disambiguation
Languages like English and Chinese contain homographs or polyphonic words whose pronunciation depends on context. Mispronouncing them can severely affect intelligibility and naturalness, yet current evaluations rarely account for this capability.
@@ 3.4 Recommendations
To promote fidelity and accuracy in TTS evaluation, we propose the following actionable recommendations, grounded in a reevaluation of current practices: (1) greater attention to perceptual validity and uncertainty in objective metrics. Interpretation of objective score differences should be approached with caution due to their nonlinear scaling, diminishing returns, domain-specific biases, and inherent uncertainty. We advocate reporting uncertainty estimates of predicted MOS, particularly under out-of-distribution scenarios. Without explicit consideration of uncertainty and prediction errors, small differences in predicted MOS should not be interpreted as genuine performance gains; (2) development of practical, discriminative, and scalable evaluation protocols. We encourage the development of improved subjective and objective evaluation metrics, as exemplified by Wang et al. (2025c) for subjective evaluation to address score saturation, reduce environmental inconsistencies, and enhance interpretability, and by Yang et al. (2025b) for objective evaluation to better correlate with human perception; (3) expanding the evaluation scope to underexplored dimensions. We advocate for a broader evaluation scope that includes dimensions such as long-form coherence, emotional expressiveness, punctuation sensitivity, and polyphonic word disambiguation.
@@ 4 Level Two: Ensuring Comparability, Standardization, and Transferability in TTS Evaluation
The second level of Responsible Evaluation builds upon the foundation of fidelity and accuracy established in the first level, arguing the importance of scientific rigor and fairness for meaningful system comparisons. Without standardized practices, even technically valid assessments fall short of supporting reliable cross-system comparisons or drawing generalizable conclusions. Current evaluation practices in TTS research remain fragmented, characterized by inconsistent methodologies, limited transparency, and poor metric transferability.
@@ 4.1 Challenges with Inconsistent Evaluation Practices
A primary challenge to comparability stems from the inconsistent usage of evaluation datasets. The most commonly used test set, LibriSpeech Panayotov et al. (2015) test-clean, is employed in divergent ways across various TTS studies. For example, VALL-E Wang et al. (2023a) utilizes 1234 utterances for zero-shot evaluation, while NatureSpeech3 Ju et al. (2024) and MaskGCT Wang et al. (2024b) employ only 40 utterance subsets, and F5-TTS Chen et al. (2025) uses 1127 utterances with punctuation and capitalization. Such disparities in test set size significantly influence evaluation metrics like WER, as detailed in Appendix A, making cross-study comparisons unreliable. Furthermore, most TTS studies do not release their prompt speech lists, while a few Wang et al. (2023a) only describe how the prompt lists are constructed. However, the sequence of prompt speech can impact performance, making results difficult to reproduce or compare.
@@
Inference tasks to evaluate zero-shot TTS are also fragmented. VALL-E Wang et al. (2023a) introduced two tasks, Continuation, which uses the first three seconds of an utterance as a prompt and continues the speech, and Cross-Sentence, which prompts with a full utterance from the same speaker. However, later work such as E2 TTS Eskimez et al. (2024) redefines the Continuation task by using the last three seconds of a truncated segment as the prompt. These inconsistencies in task definition lead to incomparable evaluation results across different works.
@@
The computation of SIM scores also varies across studies. SIM-o measures the similarity between the synthesized speech and the original prompt, while SIM-r measures the similarity between the synthesized speech and the reconstructed prompt. SIM-r is not comparable across systems using different reconstruction methods. Even for SIM-o, evaluation practices differ. VALL-E Wang et al. (2023a) excludes the prompt segment from the synthesized audio when computing similarity, whereas VALL-E 2 Chen et al. (2024) includes the prompt in both the synthesized and reference speech. As detailed in Appendix B, this leads to incomparability across different works.
@@
Widely adopted MOS evaluations frequently depart from recommended standards. While ITU-T P.808 Naderi and Cutler (2020) provides detailed protocols for conducting listening tests, many studies refer to MOS without reporting essential details, including rating scale definitions, rater calibration, playback conditions, and whether listeners rated naturalness or overall quality. Such inconsistencies reduce the reliability and comparability of MOS scores.
@@ 4.2 Challenges with Transparency in Evaluation Reporting
The reporting of MOS in TTS research also lacks transparency. Despite the importance of standardized reporting in human evaluations, many TTS studies underreport details of testing methodologies. Information regarding listener recruitment, screening procedures, compensation, and the evaluation interface is often omitted, which complicates the assessment of result replicability.
@@ 4.3 Challenges with Metric Transferability
MOS evaluations inherently lack transferability across studies. Direct comparisons of MOS scores across studies are unreliable due to the subjective nature of MOS Kirkland et al. (2023). Instead, any new comparison requires both new and previously generated speech to be jointly re-evaluated within the same subjective listening test.
@@ 4.4 Recommendations
To advance comparability, standardization, and transferability in TTS evaluation, we propose the following actionable recommendations: (1) clear distinctions between comparable and incomparable results in evaluation reporting. Metrics derived under different datasets, tasks, or configurations must not be treated as interchangeable. Any deviations should be reported explicitly to avoid misleading comparisons; (2) stick to existing standardized evaluation protocols. When formal standards such as ITU-T P.808 for MOS are available, researchers should adhere to them consistently. In the absence of formal standards, alignment with widely adopted practices is encouraged to promote practical convergence across studies; (3) ensuring transparency in evaluation reporting. Evaluation details should be disclosed, including but not limited to dataset splits, prompt lists, inference task definitions, metric configurations, human listening test procedures for MOS, and measurement setups for RTF; (4) development of transferable metrics. Model-based evaluation, including the use of LLMs as judges, offers a scalable and cost-effective alternative to human evaluation. We advocate for the creation and validation of metrics whose scores are reliably comparable across systems and studies without requiring simultaneous re-evaluation.
@@ 5 Level Three: Ensuring Ethical and Risk Oversight in TTS Evaluation
The third and most advanced level of Responsible Evaluation centers on the ethical and societal implications of TTS technology. While technical fidelity and scientific fairness form the foundation of sound evaluation, they are not sufficient to ensure TTS systems align with the public interest or broader goals of responsible AI. As TTS technology becomes increasingly realistic and pervasive, concerns arise, such as deepfakes, bias amplification, and privacy violations. These risks underscore the need to move beyond narrow technical performance and incorporate responsible AI principles, explicitly encompassing risk mitigation, fairness, transparency, accountability, and societal impact. Current evaluation practices often overlook these aspects. As a result, many TTS models are developed and deployed without adequate scrutiny of potential ethical and social consequences.
@@ 5.1 Challenges with Legal and Ethical Validity of Training Data
A core concern in responsible TTS evaluation lies in the legal and ethical status of training data. Many models are trained on large-scale speech datasets Ma et al. (2024); Chen et al. (2021a); He et al. (2024) collected from public or semi-public sources, which may lack clear copyright clearance, informed consent, or transparent data provenance. This is particularly problematic given that voice data are personally identifiable and biometric. These unresolved issues create ethical and legal blind spots. Ambiguous licensing, absence of consent for voice use or imitation, and opaque sourcing practices undermine the legitimacy of TTS systems and expose developers to reputational and legal risks. However, current evaluation practices rarely address these issues explicitly, weakening alignment with responsible AI principles in terms of data transparency, accountability in dataset construction, respect for individual consent, and protection of personal privacy.

### v2

@@ 3.1 Challenges with Objective Metrics
Objective metrics are valued for scalability and reproducibility, but they face two fundamental limitations. First, the relationship between metric scores and human perception is nonlinear and even non-monotonic, such that improvements in metric values do not necessarily translate into proportional gains in perceived quality. This discrepancy is often attributed to a mismatch between model training objectives and practical evaluation protocols (Wang et al., 2026). Second, metrics derived from neural models inevitably embody their internal biases and uncertainty (Wang et al., 2024a), rendering evaluation outcomes dependent not only on the input data but also on the metric model itself.
WER
To evaluate intelligibility and robustness, WER is computed by comparing ASR transcriptions of synthetic speech with reference texts. While effective at identifying severe intelligibility failures, its reliability is limited in two respects. First, inherent errors in ASR systems (Hsu et al., 2021; Radford et al., 2023) can lead to a mismatch between metric scores and actual perceptual quality, even when the synthesized speech is perceptually adequate to humans. Second, WER is not linearly correlated with perceived intelligibility. A low WER does not necessarily guarantee accurate transmission of key information (Tee et al., 2026). As demonstrated by the experiments in Appendix C, once WER reaches a sufficiently low level, further reductions yield negligible perceptual impact.
@@
To assess speaker similarity, the SIM score is computed by the cosine similarity between speaker embeddings extracted from reference and synthesized speech. These embeddings, derived from speaker verification models like x-vectors (Snyder et al., 2018) and ECAPA-TDNN (Desplanques et al., 2020), can be sensitive to channel variations, background noise, and even phonetic content, leading to unstable scores (Chen et al., 2022). More fundamentally, these models are trained with discriminative objectives for speaker identity classification, which are misaligned to quantify continuous perceptual similarity. In practice, once the SIM score exceeds a certain threshold, further improvements offer limited perceptual gains (Wester et al., 2016).
@@
Predicted MOS scores are generated by models trained on human ratings (Cooper and Yamagishi, 2021; Liu et al., 2025) collected following ITU-T P.808 (Naderi and Cutler, 2020). While these models offer a scalable alternative to human evaluation, they struggle with generalization and uncertainty estimation, primarily due to limitations in the diversity of training data and model representational power. Prior works (Wang et al., 2025c; Cooper et al., 2022) have shown that existing MOS prediction models often produce inconsistent results even on in-domain data, and their performance degrades significantly when applied to out-of-domain data. A typical example of domain mismatch is the widespread use of DNSMOS (Reddy et al., 2021; Cumlin et al., 2024; Reddy et al., 2022), which is trained on speech enhancement data yet commonly employed to evaluate synthesized speech. Moreover, MOS prediction models generally lack uncertainty estimation (Wang et al., 2024a), as they typically provide only point estimates without associated confidence intervals, making it difficult to assess the reliability of the predicted quality scores. This remains rarely examined in current research.
@@
To assess speech prosody, most evaluation practices (Galdino et al., 2025) employ log $F_{0}$ RMSE. However, this metric correlates weakly with human perceptual judgments (Yang et al., 2026a), and it only captures pitch but overlooks other prosodic aspects, including rhythm, stress, and intensity (Arvaniti, 2020).
@@ 3.2 Challenges with Subjective Metrics
Subjective evaluation remains the primary choice for assessing perceptual quality in TTS, with MOS serving as the dominant protocol. MOS employs a five-point absolute category rating scale to rate individual utterances. Alternative protocols such as CMOS and MUSHRA are used for pairwise or comparative assessments. Although broadly regarded as the gold standard, these methods fall short in terms of sensitivity, consistency, and practical feasibility. One major drawback of MOS stems from its limited resolution. As the quality of synthetic speech continues to improve, MOS scores tend to saturate (Wang et al., 2025e). This ceiling effect obscures perceptual differences across high-performing systems, making it increasingly difficult to distinguish among them and judgments sensitive to listener bias and preference. Another issue arises from the inherent variability in subjective ratings. Factors such as listener bias, contextual framing, playback conditions, and even day-to-day mood can introduce substantial noise. Without rigorous rater calibration and experimental controls, evaluations become unreliable. Moreover, the high cost associated with subjective evaluations presents a practical barrier. The process of recruiting a large and diverse pool of listeners, along with the need to ensure controlled testing conditions, demands considerable time and resources. These requirements often limit the feasibility and scale of such evaluations.
@@ 3.3 Underexplored Dimensions in TTS Evaluation
Existing evaluation dimensions in TTS fail to keep pace with the growing complexity of real-world applications. Widely used metrics capture only a narrow portion of what matters in practical synthesis scenarios (Manku et al., 2025). We therefore identify the following key evaluation dimensions that are essential for forward-looking assessment.
Mathematical Symbols and Formulas
Modern TTS systems like Qwen3TTS (Hu et al., 2026) are increasingly deployed in educational, scientific, and accessibility-oriented scenarios, where accurate verbalization of mathematical symbols, formulas, and structured notations is critical. Mathematical expressions often exhibit non-linear and deeply nested structures, implicit grouping, and context-dependent reading conventions that are poorly handled by current text normalization pipelines. Errors in symbol pronunciation, operator scope, or structural cues can severely impair comprehension, yet often remain invisible to ASR-based evaluations. Beyond mathematics, real-world scenarios often interleave formulas with diverse content, further complicating evaluation. While recent efforts such as EmergentTTS-Eval (Manku et al., 2025) begin to cover emails, phone numbers, URLs, addresses, STEM equations, units, and notations, the community still lacks multi-domain benchmarks and evaluation protocols that systematically assess symbolic and structured speech synthesis, leaving a gap between real-world requirements and current evaluation practices.
@@
In real-world applications such as audiobooks and podcasts, coherence across sentences and stability in prosody and speaker identity are essential. However, most existing evaluations center on short utterances such as LibriTTS (Zen et al., 2019) and Seed-TTS-eval (Anastassiou et al., 2024). There is a lack of representative test sets and metrics specifically designed to assess long-form fluency, consistency, and discourse-level control.
@@
Recent TTS models have demonstrated increasing capability in synthesizing expressive speech (Du et al., 2024b; Hu et al., 2026), yet evaluation methodologies remain underdeveloped. In particular, there is no consensus on emotion taxonomies or scales for emotion intensity, and subjective metrics like emotion MOS often lack sensitivity to subtle distinctions (Yang et al., 2025a). Moreover, widely used emotional speech datasets (Busso et al., 2008; Livingstone and Russo, 2018; Cao et al., 2014) primarily rely on discrete labels and provide limited coverage of expressive diversity.
@@ 3.4 Recommendations
To promote fidelity and accuracy in TTS evaluation, we propose the following actionable recommendations, grounded in a reevaluation of current practices: (1) greater attention to perceptual validity and uncertainty in objective metrics. Interpretation of objective score differences should be approached with caution due to their nonlinear scaling, diminishing returns, domain-specific biases, and inherent uncertainty. We advocate reporting uncertainty estimates of predicted MOS, particularly under out-of-distribution scenarios. Without explicit consideration of uncertainty and prediction errors, small differences in predicted MOS should not be interpreted as genuine performance gains; (2) development of practical, discriminative, and scalable evaluation protocols. We encourage the development of improved subjective and objective evaluation metrics, as exemplified by (Wang et al., 2025e) for subjective evaluation, which addresses score saturation, reduces environmental inconsistencies, and enhances interpretability, and by (Tee et al., 2026), which evaluates the accuracy of key information in synthesized speech beyond the word level. (3) expanding the evaluation scope to underexplored dimensions. We advocate for a broader evaluation scope that includes dimensions such as long-form coherence, emotional expressiveness, and punctuation sensitivity.
@@ 4 Level Two: Ensuring Comparability, Standardization, and Transferability in TTS Evaluation
The second level of Responsible Evaluation builds upon the foundation of fidelity and accuracy established in the first level, arguing the importance of scientific rigor to enable reliable and fair cross-system comparisons. Without standardized practices, even technically valid assessments fall to support meaningful comparison or generalizable conclusions. Current evaluation practices in TTS research remain fragmented, characterized by inconsistent methodologies, limited transparency, and poor metric transferability.
@@ 4.1 Challenges with Inconsistent Evaluation Practices
A primary challenge to comparability stems from the inconsistent usage of evaluation datasets. The most commonly used test set, LibriSpeech (Panayotov et al., 2015) test-clean, is employed in divergent ways across various TTS studies. For example, VALL-E (Wang et al., 2023) utilizes 1234 utterances for zero-shot evaluation, while NatureSpeech3 (Ju et al., 2024) and MaskGCT (Wang et al., 2024c) employ only 40 utterance subsets, and F5-TTS (Chen et al., 2025) uses 1127 utterances with punctuation and capitalization. Such disparities in test set size significantly influence evaluation metrics like WER, as detailed in Appendix A, making cross-study comparisons unreliable. Moreover, most TTS studies do not release prompt speech lists. However, the sequence of prompt speech can impact performance, making results difficult to reproduce.
@@
Inference tasks to evaluate zero-shot TTS are also fragmented. VALL-E (Wang et al., 2023) introduced two tasks, Continuation, which uses the first three seconds of an utterance as a prompt and continues the speech, and Cross-Sentence, which prompts with a full utterance from the same speaker. However, later work such as E2 TTS (Eskimez et al., 2024) redefines the Continuation task by using the last three seconds of a truncated segment as the prompt. These inconsistencies in task definition lead to incomparable evaluation results across different works.
@@
The computation of SIM scores varies across studies. SIM-o measures the similarity between the synthesized speech and the original prompt, while SIM-r measures the similarity between the synthesized speech and the reconstructed prompt. SIM-r is not comparable across systems using different reconstruction methods. Evaluation practices for SIM-o also differ: VALL-E (Wang et al., 2023) excludes the prompt segment when computing similarity, while VALL-E 2 (Chen et al., 2024) includes the prompt. As detailed in Appendix B, these differences lead to incomparability across works.
@@
Widely adopted MOS evaluations frequently depart from recommended standards. While ITU-T P.808 (Naderi and Cutler, 2020) provides detailed protocols for conducting listening tests, many studies refer to MOS without reporting essential details, including rating scale definitions, rater calibration, playback conditions, and whether listeners rated naturalness or overall quality. Such inconsistencies reduce the reliability and comparability of MOS scores.
@@ 4.2 Challenges with Transparency in Evaluation Reporting
The reporting of MOS in TTS research often lacks transparency (Wang et al., 2025e). Despite the importance of standardized reporting in human evaluations, many TTS studies underreport details of testing methodologies. Information regarding listener recruitment, screening procedures, compensation, and the evaluation interface is often omitted, which complicates the assessment of result replicability.
@@ 4.3 Challenges with Metric Transferability
MOS evaluations are not transferable across studies. Direct comparisons of MOS scores across studies are meaningless due to the subjective nature of MOS (Kirkland et al., 2023). Instead, any new comparison requires both new and previously generated speech to be jointly re-evaluated within the same subjective listening test.
@@ 4.4 Recommendations
To advance comparability, standardization, and transferability in TTS evaluation, we propose the following actionable recommendations: (1) clear distinctions between comparable and incomparable results in evaluation reporting. Metrics derived from different datasets, tasks, or configurations must not be treated as interchangeable. Any deviations should be reported explicitly to avoid misleading comparisons. (2) stick to existing standardized evaluation protocols. When formal standards such as ITU-T P.808 for MOS are available, researchers should adhere to them consistently. In the absence of formal standards, alignment with widely adopted practices is encouraged to promote practical convergence across studies. (3) ensuring transparency in evaluation reporting. Evaluation details should be disclosed, including but not limited to dataset splits, prompt lists, inference task definitions, metric configurations, human listening test procedures for MOS, and measurement setups for RTF. (4) development of transferable metrics. Model-based evaluation, including recent LLMs-as-judges, offers a scalable alternative to human evaluation. We encourage the development of objective metrics that better align with human perception and yield scores that are reliably comparable across systems without requiring re-evaluation, as exemplified by (Yang et al., 2026a) for prosody diversity and (Yang et al., 2026b) for style similarity.
@@ 5 Level Three: Ensuring Ethical and Risk Oversight in TTS Evaluation
The third and most advanced level of Responsible Evaluation centers on the ethical and societal implications of TTS technology. While technical fidelity and scientific fairness form the foundation of sound evaluation, they are not sufficient to ensure TTS systems align with the public interest or broader goals of responsible AI. As TTS technology becomes increasingly realistic and pervasive, concerns arise, such as deepfakes, bias amplification, and privacy violations. These risks underscore the need to move beyond narrow technical performance and incorporate responsible AI principles, explicitly encompassing risk mitigation, fairness, transparency, accountability, and societal impact. Current evaluation practices often overlook these aspects. As a result, many TTS systems are developed or deployed without adequate scrutiny of potential ethical and social consequences.
@@ 5.1 Challenges with Legal and Ethical Validity of Training Data
A core concern in responsible TTS evaluation lies in the legal and ethical status of training data. Many models are trained on large-scale speech datasets (Ma et al., 2024; Chen et al., 2021a; He et al., 2024) collected from public or semi-public sources, which may lack clear copyright clearance, informed consent, or transparent data provenance. This is particularly problematic given that voice data are personally identifiable and biometric. These unresolved issues create ethical and legal blind spots. Ambiguous licensing, absence of consent for voice use or imitation, and opaque sourcing practices undermine the legitimacy of TTS systems and expose developers to reputational and legal risks. However, current evaluation practices rarely address these issues explicitly, weakening alignment with responsible AI principles in terms of data transparency, accountability in dataset construction, respect for individual consent, and protection of personal privacy.